SOHO Daily Meeting Minutes for Tuesday 24 June 2003

Status Report From: Goddard Space Flight Center
Posted: Tuesday, June 24, 2003


* We expect to lose the high gain antenna link on Wednesday or Thursday of this week for ~20 days. Starting Wednesday, we are scheduling daily passes on 34 meter stations for monitoring and commanding of the spacecraft. The 34 meter schedule for this week is included below. Since the times are very limited, we'd like to request that ALL TEAMS BE AVAILABLE (at EOF or by phone) during the passes, in case your instrument comes up in an unexpected state. When downlink only 34 meter stations are scheduled, a 26 meter station will be used for uplink (i.e. all 34 meter passes will have commanding capability).

Week 26 34-Meter Passes

                   BOT   EOT
   Wed     DSS-65 15:05-18:25 UT D/L only and DSS-16 26-meter overlap        
   Thurs   DSS-54 15:30-17:45 UT
   Fri     DSS-54 13:00-17:45 UT
   Sat     DSS-24 17:35-20:35 UT
   Sun     DSS-54 13:00-17:30 UT

* Current RHESSI default target is AR10386. When selecting this target to maximize coordination w/others, use campaign number 6850.

* EIT Shutterless Segment #2 is today. No NRT will be available from 19-21:40 UT.

* GITT scans for the EOF and EAF are scheduled for Thursday 26 June at 9:30am.


Spacecraft Status: Nominal, except for the HGA pointing mechanism HGA Position Monitoring Disabled HGA Nominal Control Disabled

Spacecraft Anomalies: None

Accomplished Activities: CELIAS, SWAN, NONCOHO_03174 OCD #1799: Z-Axis High Speed Movement with APME-A, Stable Motor Temps

Planned Activities: RSL

New tracking star, mag = 2.9

Upcoming Operations: None

Ground Anomalies:

173/1611 Wayward echo block, no operational impacts.

174/1616 Received degraded TLM from 1616 to 1712 due to a line overflow at JPL caused by a VLBI test being performed.


06/23 DOY 174 - 06/24 DOY 175

EIT daily On-Line Solar Status and observations:

LASCO daily On-Line Solar Status and observations:

Web Page for Planning:

Magnetic Maps and Heliospheric Forecasts:


CDS: Safed at 12 UT today

UVCS: Nominal. JOP155 until 16 UT and then, safing at ~17:30 UT

LASCO: Doors closed since yesterday at 18:30 UT

EIT: Nominal. 195 CME watch and synoptics, shutterless, then bakeout.

MDI: Nominal. Full disk magnetograms and dopplergrams, EIT shutterless support then safing at 22 UT

TRACE: Nominal. JOP155 and EIT shutterless support on AR10386 and its surrounding region.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.